# KungfuBot: Physics-Based Humanoid Whole-Body Control for Learning Highly-Dynamic Skills

### v1

@@ 1 Introduction
Recently, several RL-based whole-body control frameworks have been proposed to track motions expressive ; humanplus , which often take a reference kinematic motion as input and output the control actions for a humanoid robot to imitate it. To address physical feasibility issues, H2O and OmniH2O he2024h20 ; OmniH2O remove the infeasible motions using a trained privileged imitation policy, producing a clean motion dataset. ExBody expressive constructs a feasible motion dataset by filtering via language labels, such as ‘wave’ and ‘walk’. Exbody2 exbody2 trains an initial policy on all motions and uses the tracking error to measure the difficulty of each motion. However, it would be costly to train the initial policy and find an optimal dataset. There is also a lack of suitable tolerance mechanisms for difficult-to-track motions in the training process. As a result, previous methods are only capable of tracking low-speed and smooth motions. Recently, ASAP he2025asap introduces a multi-stage mechanism and learned a residual policy to compensate for the sim-to-real gap, reducing the difficulties in tracking agile motions. However, ASAP involves a total of four training stages, and the training of residual policy requires MoCap systems to record real-robot states.
@@ 2 Preliminaries
Problem Formulation. We adopt the Unitree G1 robot unitree-g1 in our work, which has 23 degrees of freedom (DoFs) to control, excluding the 3 DoFs in each wrist of the hand. We formulate the motion imitation problem as a goal-conditional RL problem with Markov Decision Process $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{S}^{\rm ref},\gamma,r,P)$, where $\mathcal{S}$ and $\mathcal{S}^{\rm{ref}}$ are the state spaces of the humanoid robot and reference motion, respectively, $\mathcal{A}$ is the robot’s action space, $r$ is a mixed reward function consisting motion-tracking and regularization rewards, and $P$ is the transition function depending on the robot morphology and physical constraints. At each time step $t$, the policy $\pi$ observes the proprioceptive state $\bm{s}_{t}^{\rm{prop}}$ of the robot and generates action $\bm{a}_{t}$, with the aim of obtaining the next-state $\bm{s}_{t+1}$ that follows the corresponding reference state $\bm{s}^{\rm ref}_{t+1}$ in the reference trajectory $[\bm{s}^{\rm{ref}}_{0},\ldots,\bm{s}^{\rm ref}_{N-1}]$. The action $\bm{a}_{t}\in\mathbb{R}^{23}$ is the target joint position for a PD controller to compute the motor torques. We adopt an off-the-shelf RL algorithm, PPO ppo , for policy optimization with an actor-critic architecture.
@@ 3.2.1 Exponential Form Tracking Reward
The reward function in PBHC, detailed in Appendix C.2, comprises two components: task-specific rewards, which enforce accurate tracking of reference motions, and regularization rewards, which promote overall stability and smoothness.
@@ 3.2.2 Optimal Tracking Factor
[…]
@@ 5.3 Main Result
As shown in Table 1, PBHC consistently outperforms the baselines OmniH2O and ExBody2 across all evaluation metrics. These improvements can be attributed to our adaptive motion tracking mechanism, which automatically adjusts tracking factors based on motion characteristics, whereas the fixed, empirically tuned parameters in the baselines fail to generalize across diverse motions. While MaskedMimic performs well on certain metrics, it is primarily designed for character animation and is not deployable for robot control, as it does not account for constraints such as partial observability and action smoothness. Therefore, we treat it as an oracle-style lower bound rather than a directly comparable baseline.
Table 1: Main results comparing different methods across difficulty levels. PBHC consistently outperforms deployable baselines and approaches oracle-level performance. Results are reported as mean $\pm$ one standard deviation. Bold indicates methods within one standard deviation of the best result, excluding MaskedMimic.
@@ 5.5 Real-World Deployment
As shown in Fig. 8, 11 and the supporting videos, our robot in real world demonstrates outstanding dynamic capabilities through a diverse repertoire of advanced skills: (1) sophisticated martial arts techniques including powerful boxing combinations (jabs, hooks, and horse-stance punches) and high-degree kicking maneuvers (front kicks, jump kicks, side kicks, back kicks, and spinning roundhouse kicks); (2) acrobatic movements such as full 360-degree spins; (3) flexible motions including deep squats and stretches; (4) artistic performances ranging from dynamic dance routines to graceful Tai Chi sequences. This comprehensive skill set highlights our system’s remarkable versatility, dynamic control, and real-world applicability across both athletic and artistic domains.
@@
MuJoCo 33.18$\pm$2.720 1061.24$\pm$83.27 2.96$\pm$0.342 2.90$\pm$0.498 67.71$\pm$6.747
Real 36.64$\pm$2.592 1130.05$\pm$9.478 3.01$\pm$0.126 3.12$\pm$0.056 65.68$\pm$1.972
@@ 6 Conclusion & Limitations
This work is supported by the National Natural Science Foundation of China (Grant No.62306242), the Young Elite Scientists Sponsorship Program by CAST (Grant No. 2024QNRC001), and the Yangfan Project of the Shanghai (Grant No.23YF11462200).

### v2

@@ 1 Introduction
Recently, several RL-based whole-body control frameworks have been proposed to track motions expressive ; humanplus , which often take a reference kinematic motion as input and output the control actions for a humanoid robot to imitate it. To address physical feasibility issues, H2O and OmniH2O he2024h20 ; OmniH2O remove the infeasible motions using a trained privileged imitation policy, producing a clean motion dataset. ExBody expressive constructs a feasible motion dataset by filtering via language labels, such as ‘wave’ and ‘walk’. Exbody2 exbody2 trains an initial policy on all motions and uses the tracking error to measure the difficulty of each motion. However, it would be costly to train the initial policy and find an optimal dataset. There is also a lack of suitable tolerance mechanisms for difficult-to-track motions in the training process. As a result, previous methods are only capable of tracking low-speed and smooth motions. Recently, ASAP he2025asap introduces a multi-stage mechanism and learned a residual policy to compensate for the sim-to-real gap, reducing the difficulties in tracking agile motions. Unlike ASAP, we focus on improving motion feasibility and agility entirely in simulation.
@@ 2 Preliminaries
Problem Formulation. We adopt the Unitree G1 robot unitree-g1 in our work, which has 23 degrees of freedom (DoFs) to control, excluding the 3 DoFs in each wrist of the hand. We formulate the motion imitation problem as a goal-conditional RL problem with Markov Decision Process $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{S}^{\rm ref},\gamma,r,P)$, where $\mathcal{S}$ and $\mathcal{S}^{\rm{ref}}$ are the state spaces of the humanoid robot and reference motion, respectively, $\mathcal{A}$ is the robot’s action space, $r$ is a mixed reward function consisting motion-tracking and regularization rewards, and $P$ is the transition function depending on the robot morphology and physical constraints. At each time step $t$, the policy $\pi$ observes the proprioceptive state $\bm{s}_{t}^{\rm{prop}}$ of the robot and generates action $\bm{a}_{t}$, with the aim of obtaining the next-state $\bm{s}_{t+1}$ that follows the corresponding reference state $\bm{s}^{\rm ref}_{t+1}$ in the reference trajectory $\bm{s}^{\rm{ref}}_{0},\ldots,\bm{s}^{\rm ref}_{N-1}]$. The action $\bm{a}_{t}\in\mathbb{R}^{23}$ is the target joint position for a PD controller to compute the motor torques. We adopt an off-the-shelf RL algorithm, PPO ppo , for policy optimization with an actor-critic architecture.
@@ 3.2.1 Exponential Form Tracking Reward
The reward function in PBHC, detailed in Appendix C.2, comprises two components: task‐specific rewards, which enforce accurate tracking of reference motions, and regularization rewards, which promote overall stability and smoothness.
@@ 3.2.2 Optimal Tracking Factor
$\max_{\sigma\in\mathbb{R_{+}}}$ ${J}^{\mathrm{ex}}(\bm{x}^{*}),\qquad\text{s.t.}\>\>$ $\bm{x}^{*}\in\arg\max_{\bm{x}\in\mathbb{R}^{N}_{+}}J^{\mathrm{in}}(\bm{x},\sigma)+R(\bm{x}).$ (6)
This simplified modeling provides an intuitive connection to the RL training process.
  • The lower-level optimization represents the standard RL procedure, where a policy is trained to maximize tracking reward and other reward terms, given a specific $\sigma$.
  • The upper-level optimization, outside the RL loop, selects $\sigma$ to minimize the total tracking error of the final converged policy. This outer optimization is not reward maximization but a performance-driven objective based on absolute external metrics.
@@ 5.3 Main Result
As shown in Table 1, PBHC consistently outperforms the baselines OmniH2O and ExBody2 across all evaluation metrics. These improvements can be attributed to our adaptive motion tracking mechanism, which automatically adjusts tracking factors based on motion characteristics, whereas the fixed, empirically tuned parameters in the baselines fail to generalize across diverse motions. While MaskedMimic performs well on certain metrics, it is primarily designed for character animation and is not deployable for robot control, as it does not account for constraints such as partial observability and action smoothness. To enable a fair comparison, we also train an oracle version of PBHC that similarly overlooks such constraints, in the same manner as MaskedMimic.
Table 1: Main results comparing different methods across difficulty levels. PBHC consistently outperforms deployable baselines and approaches oracle-level performance. Results are reported as mean $\pm$ one standard deviation. Bold indicates methods within one standard deviation of the best result, excluding Oracle baselines. Asterisks (*) denote significant improvements ($p<0.05$) of our method over baselines per two-sided permutation tests.
@@ 5.5 Real-World Deployment
To investigate Q4, we deploy the policies in real robot. As shown in Fig. 8, 12 and the supporting videos, our robot in real world demonstrates outstanding dynamic capabilities through a diverse repertoire of advanced skills: (1) sophisticated martial arts techniques including powerful boxing combinations (jabs, hooks, and horse-stance punches) and high-degree kicking maneuvers (front kicks, jump kicks, side kicks, back kicks, and spinning roundhouse kicks); (2) acrobatic movements such as full 360-degree spins; (3) flexible motions including deep squats and stretches; (4) artistic performances ranging from dynamic dance routines to graceful Tai Chi sequences. This comprehensive skill set highlights our system’s remarkable versatility, dynamic control, and real-world applicability across both athletic and artistic domains.
@@
[…]
5.6 Learning Curves
To additionally illustrate the training process and verify its stability, we present in Fig. 9 the learning curves for three representative motions—Jabs Punch, Tai Chi, and Roundhouse Kick—showing both the mean episode length and mean reward. These curves provide an intuitive view of how the policy improves over time, and it can be observed that training gradually stabilizes and converges after approximately 20k steps, demonstrating the reliability and efficiency of our approach in learning complex motion behaviors.
Figure 9: Mean episode length and mean reward across three motions. Both curves indicate that training gradually stabilizes after 20k steps.
@@ 6 Conclusion & Limitations
This work is supported by the National Key Research and Development Program of China (Grant No.2024YFE0210900), Shanghai Municipal Science and Technology Major Project (Grant No.2021SHZDZX0102), the National Natural Science Foundation of China (Grant No.62306242 and No.62322603), the Young Elite Scientists Sponsorship Program by CAST (Grant No.2024QNRC001), and the Yangfan Project of the Shanghai (Grant No.23YF11462200).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The paper’s contributions and scope are specifically claimed in the abstract and introduction. Our key contributions are summarized in Section 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitation of our work in Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We propose and proof our theory in Section 3 and provide more detailed analysis in Appendix A.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We fully describe our settings in simulation and real world experiments in Section 5, Appendix B, Appendix C and Appendix D, and the information provided is enough to the reproducibility of our main experimental results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide sufficient material in supplemental material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specify our training and test details in Section 5, Appendix C and Appendix D.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We define the evaluation metrics in Section 5.1. And then we compare our method with other baselines, showing statistical results in Section 5.3.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Please see our experimental details in Appendix D.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We read the NeurIPS Code of Ethics, and make sure that our research conducted in the paper conform with the NeurIPS Code of Ethics in every respect.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the societal impacts in Appendix E.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: No such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite works of related assets.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: No new assets are introduced.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: NA
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: NA
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: NA
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.